OpenStudy (anonymous):

Find the standard form for the equation of a circle (x−h)2+(y−k)2=r2 with a diameter that has endpoints (−8,6) and (9,−8)

OpenStudy (mertsj):

Use the distance formula to find the diameter.

OpenStudy (mertsj):

Then use the midpoint formula to find the center of the circle.

OpenStudy (anonymous):

sqrt((9-(-8))^2(((-8)-6)^2)=238

OpenStudy (anonymous):

(1/2,-1)=midpoint

OpenStudy (mertsj):

\[\sqrt{(x _{2}-x _{1})^2+(y _{2}-y _{1})^2}=d\]

OpenStudy (mertsj):

The midpoint is correct.

OpenStudy (anonymous):

what about the distance formula

OpenStudy (mertsj):

\[\sqrt{(-8-9)^2+(6+8)^2}=d=\sqrt{289+196}=\sqrt{485}=r\]

OpenStudy (mertsj):

So r^2=485

OpenStudy (mertsj):

And now you can write the equation of the circle because you know h, k and r^2

OpenStudy (anonymous):

oh ok i see what i did wrong thank you

OpenStudy (mertsj):

yw